# Lie algebras generated by reflections in types BCD

### v1

@@ 4.3.3. Arm and leg (A&L) bipartitions
This is seen via the same line of reasoning as given in the proof of [13, Lemme 12], applied now to the natural representation $S^{\beta}$ for $\mathcal{B}_{n}$. In this case one must also consider the action of the generator $t_{1}$, for which one checks that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(t_{1})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(t_{1})+(d-1)\cdot\operatorname{Id}$. This is the same relation as for the generators $s_{1},\ldots,s_{n-1}$, so the reasoning in the proof of [13, Lemme 12] applies. ∎
@@
Let $\rho_{\operatorname{Grp}}^{\wedge^{d}}:\mathfrak{b}_{n}\to\operatorname{End}(\Lambda^{d}(S^{\beta}))$ and $\rho_{\operatorname{Lie}}^{\wedge^{d}}:\mathfrak{b}_{n}\to\operatorname{End}(\Lambda^{d}(S^{\beta}))$ be the $\mathfrak{b}_{n}$-module structure maps for $\operatorname{Res}^{\mathbb{C}\mathcal{B}_{n}}_{\mathfrak{b}_{n}}(\Lambda^{d}(S^{\beta})\otimes\varepsilon^{\prime})$ and $\Lambda^{d}(\operatorname{Res}^{\mathbb{C}\mathcal{B}_{n}}_{\mathfrak{b}_{n}}(S^{\beta}\otimes\varepsilon^{\prime}))$, respectively. The proof again mirrors that of [13, Lemme 12]. In this case one sees that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(t_{1})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(t_{1})-(d-1)\cdot\operatorname{Id}$, while for $s_{1},\ldots,s_{n-1}$ one again gets $\rho_{\operatorname{Lie}}^{\wedge^{d}}(s_{i})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(s_{i})+(d-1)\cdot\operatorname{Id}$. Thus, for each element $s$ in a set of generators for $\mathfrak{b}_{n}$, one sees that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(s)$ and $\rho_{\operatorname{Grp}}^{\wedge^{d}}(s)$ differ by a central element of $\operatorname{End}(\Lambda^{d}(S^{\beta}))$. This implies that the restrictions of $\rho_{\operatorname{Lie}}^{\wedge^{d}}$ and $\rho_{\operatorname{Grp}}^{\wedge^{d}}$ to $\mathfrak{b}_{n}^{\prime}$ are the same. ∎
@@ 4.3.4. Improper bipartitions
The structure of $\mathfrak{s}_{\lambda}$ was determined by Marin in [13]. We summarize his results. Say that a partition $\lambda$ is a hook if $\lambda=[n-d,1^{d}]$ for some $10\leq d\leq n-$ 1.333Marin refers to partitions of this form as équerres. Let
$E_{n}$ =$=\left\{\lambda\vdash n:\text{$\lambda$ is not a hook and $\lambda\neq\lambda^{*}$}\right\},$
$F_{n}$ =$=\left\{\lambda\vdash n:\text{$\lambda$ is not a hook and $\lambda=\lambda^{*}$}\right\}.$
@@
  • Given $\lambda\in F_{n}$, let $b(\lambda)$ denote the length of the diagonal in the Young diagram of $\lambda$. Then there exists a nondegenerate bilinear form [^4] Up to a nonzero scalar multiple, this is the form (|){$(-|-)_{\left\{\lambda,\emptyset\right\}}$ on $S^{\left\{\lambda,\emptyset\right\}}=S^{\lambda}$ defined prior to Lemma 4.3.2. (|)(\;|\;)( | ) on $S^{\lambda}$, which is symmetric (resp. anti-symmetric) if $(n-b(\lambda))/2$ is even (resp. odd), and with respect to which
@@
  • For $10\leq d\leq n-$ 1, set $\alpha_{d}=\alpha_{n,d}=[n-d,1^{d}]$, and set $\alpha=\alpha_{n,1}$. Then $S^{\alpha_{d}}\cong\Lambda^{d}(S^{\alpha})$ as $\mathfrak{s}_{n}^{\prime}$-modules by [13, Lemme 12]. The map $\rho_{\alpha_{d}}:\mathfrak{s}_{n}^{\prime}\to\mathfrak{sl}(\alpha_{d})$ factors through $\rho_{\alpha}:\mathfrak{s}_{n}^{\prime}\to\mathfrak{sl}(\alpha)$, and $\mathfrak{s}_{\alpha}=\mathfrak{sl}(\alpha)\cong\mathfrak{sl}(n-1)$. Then $\mathfrak{s}_{\alpha_{d}}\cong\mathfrak{s}_{\alpha}$ if $21\leq d\leq n-$ 2, and $\mathfrak{s}_{\alpha_{d}}=0$ otherwise.
Theorem 4.3.10 ([13, Theorem A]).
@@ 4.3.5. Summary
$E(n)$ =$=\left\{(\lambda,\mu)\in\mathcal{BP}(n)^{\circ}:\text{$(\lambda,\mu)$ is not an A\&L and $(\lambda,\mu)\neq(\mu^{*},\lambda^{*})$}\right\},$
$F(n)$ =$=\left\{(\lambda,\mu)\in\mathcal{BP}(n)^{\circ}:\text{$(\lambda,\mu)$ is not an A\&L and $(\lambda,\mu)=(\mu^{*},\lambda^{*})$}\right\},$
$E\{n\}$ =$=\left\{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}^{\circ}:\text{$\left\{\lambda,\mu\right\}$ is not an A\&L and $\left\{\lambda,\mu\right\}\neq\left\{\lambda^{*},\mu^{*}\right\}$}\right\},$
$F\{n\}$ =$=\left\{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}^{\circ}:\text{$\left\{\lambda,\mu\right\}$ is not an A\&L and $\left\{\lambda,\mu\right\}=\left\{\lambda^{*},\mu^{*}\right\}$}\right\}.$
@@ 5.1. Statement of the main theorem
Verified by dimension comparison using GAP [6]. ∎
For the induction step we let $n\geq 6$ be fixed and we assume that Theorem 5.1.1 is true for the value $n-1$. Under this assumption we first show in Sections 5.25.9 that parts (2)–(9) of Theorem 5.1.1 are true for the value $n$; part (1) is true for all $n\geq 2$ by the identifications in (4.3.14) and by Theorem 4.3.10. Then in Section 5.10 we show that the maps (4.3.16) and (4.3.17) are surjections. Since the maps are injections by construction, this finishes the proof.
In broad strokes, the proofs of parts (2)–(9) are all the same: In addition to the given unknown Lie algebra $\mathfrak{g}$ (either $\mathfrak{d}_{\beta}$, $\mathfrak{b}_{\beta}$, $\mathfrak{b}_{\gamma}$, $\mathfrak{b}_{(\lambda,\mu)}$, etc.), we consider a semisimple subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ arising from the corresponding Weyl group of rank $n-1$. From the restriction formulas in Sections 2.5 and 3.5, we first deduce that $\mathfrak{g}$ is simple using [13, Lemme 15]. Then in most cases, we see with the help of [13, Lemme 13] and [13, Lemme 14] that the rank of $\mathfrak{h}$ is sufficiently large to force $\mathfrak{g}$ to have the claimed value. There is however one case where this strategy is not entirely sufficient to determine $\mathfrak{g}$: For part (7) we must apply more detailed information about the module $S^{\left\{\lambda,\pm\right\}}$ to rule out the possibility that $\mathfrak{g}$ is either an orthogonal or symplectic Lie algebra.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (2) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The dimensions of the modules are $\dim(S^{\beta})=n$, $\dim(S^{\left\{[n-2],[1]\right\}})=n-1$, and $\dim(S^{\left\{[n-1],\emptyset\right\}})=1$. Let $\mathfrak{h}=\rho_{\beta}(\mathfrak{d}_{n-1}^{\prime})$. Then $\mathfrak{h}\subseteq\mathfrak{d}_{\beta}$ are two semisimple subalgebras of $\mathfrak{sl}(\beta)$, and from the induction hypothesis one gets $\mathfrak{h}=\mathfrak{sl}\{[n-2],[1]\}$. In particular, $\mathfrak{h}$ is semisimple of rank $\operatorname{rk}(\mathfrak{h})=n-2$, while $\operatorname{rk}(\mathfrak{d}_{\beta})\leq\operatorname{rk}(\mathfrak{sl}(\beta))=n-1$, so $\operatorname{rk}(\mathfrak{d}_{\beta})<2\cdot\operatorname{rk}(\mathfrak{h})$. Then $\mathfrak{d}_{\beta}$ is simple by [13, Lemme 15]. Now $2\cdot\operatorname{rk}(\mathfrak{d}_{\beta})\geq 2\cdot\operatorname{rk}(\mathfrak{h})=2(n-2)>n=\dim(\beta)$, so $\mathfrak{d}_{\beta}=\mathfrak{sl}(\beta)$ by [13, Lemme 13].
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (3) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.3.1 that the right side of (5.3.1) is a summand of the right side of (4.3.16) for the value $n-1$. This implies by induction that (5.3.1) is a surjection and $\mathfrak{r}_{(\nu,\lambda^{*})}\cong\mathfrak{sl}(\nu,\lambda^{*})$ for each $\nu\prec\lambda$. By the discussion of Section 4.3.1, $S^{(\nu,\lambda^{*})}$ and $S^{(\lambda,\nu^{*})}$ are dual as $\mathfrak{r}^{\prime}$-modules. Since $\dim(S^{(\nu,\lambda^{*})})\geq n-2\geq 4$ by Lemmas 5.3.1 and 2.2.1, and since the natural module for $\mathfrak{sl}(k)$ is self-dual only if $k=2$, it follows that $S^{(\nu,\lambda^{*})}\not\cong S^{(\lambda,\nu^{*})}$ as $\mathfrak{r}^{\prime}$-modules.
@@
Then $\mathfrak{b}_{(\lambda,\lambda^{*})}$ is a simple Lie algebra by [13, Lemme 15]. Since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})$, the preceding inequalities also imply that $2\cdot\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})\leq\dim(S^{(\lambda,\lambda^{*})})<4\cdot\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})$. Now since $\dim(S^{(\lambda,\lambda^{*})})\geq\binom{n}{n/2}$, it follows from [13, Lemme 14] that $\mathfrak{b}_{(\lambda,\lambda^{*})}=\mathfrak{osp}(\lambda,\lambda^{*})$. ∎
@@
    for some $21\leq i\leq n-$ 2, and $(\nu,\tau)$ is the unique arm and leg such that $(\nu,\tau)\prec(\lambda,\mu)$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (4) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.4.1 that the right side of (5.4.2) is a summand of the right side of (4.3.16) for the value $n-1$, except perhaps that a summand of the form $\mathfrak{r}_{\beta_{n-1,1}}$ or $\mathfrak{r}_{\gamma_{n-1,1}}$ may be replaced by an isomorphic summand of the form $\mathfrak{r}_{\beta_{n-1,d}}$ or $\mathfrak{r}_{\gamma_{n-1,d}}$ for some $21\leq d\leq n-$ 2, and a summand of the form $\mathfrak{s}_{\alpha_{n-1,1}}$ may be replaced by an isomorphic summand $\mathfrak{s}_{\alpha_{n-1,d}}$ for some $31\leq d\leq n-$ 3. This implies by the induction hypothesis that (5.4.2) is a surjection, and implies for each $(\nu,\tau)\prec(\lambda,\mu)$ that $\mathfrak{r}_{(\nu,\tau)}$ is a simple Lie algebra acting irreducibly on $S^{(\nu,\tau)}$.
Let $\mathfrak{h}=\rho_{(\lambda,\mu)}(\mathfrak{r}^{\prime})$. Then $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{b}_{(\lambda,\mu)}$ such that the restriction of $S^{(\lambda,\mu)}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. Since $\mathfrak{h}\subseteq\mathfrak{b}_{(\lambda,\mu)}\subseteq\mathfrak{sl}(\lambda,\mu)$, then $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{b}_{(\lambda,\mu)})\leq\dim(S^{(\lambda,\mu)})$. Our goal now, as in the proof of Lemma 5.2.1, is to show that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{(\lambda,\mu)})$. This will imply first by [13, Lemme 15] that $\mathfrak{b}_{(\lambda,\mu)}$ is a simple Lie algebra, and then by [13, Lemme 13] that $\mathfrak{b}_{(\lambda,\mu)}=\mathfrak{sl}(\lambda,\mu)$.
@@
Suppose $\mu=[1]$ and $\lambda$ is a hook partition. Then $\lambda=\alpha_{n-1,d}=[n-1-d,1^{d}]$ for some $31\leq d\leq n-$ 3; the values $d=0$ and $d=n-2$ are excluded by the assumption that $(\lambda,\mu)$ is not an arm and leg. In this case, (5.4.1) becomes
@@
From Section 4.3.4, $\mathfrak{r}_{(\alpha_{n-1,d},\emptyset)}\cong\mathfrak{sl}(\alpha_{n-1,1})\cong\mathfrak{sl}(n-2)$. For $d\neq n-3$, the bipartition $(\alpha_{n-2,d},[1])$ is not an arm and leg, and hence $\mathfrak{r}_{(\alpha_{n-2,d},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d},[1])})$, while for $d=n-3$, one has $(\alpha_{n-2,d},[1])=([1^{n-2}],[1])=\gamma_{n-1,n-2}$. By induction, $\mathfrak{r}_{\gamma_{n-1,n-2}}\cong\mathfrak{sl}(\gamma_{n-1,1})\cong\mathfrak{sl}(n-1)$, so by dimension comparison the isomorphism $\mathfrak{r}_{(\alpha_{n-2,d},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d},[1])})$ is true for $d=n-3$. Similarly, one sees for all $31\leq d\leq n-$ 3 that $\mathfrak{r}_{(\alpha_{n-2,d-1},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d-1},[1])})$. Then
@@
$0\textstyle 2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=(n-2)\cdot\binom{n-2}{d}+(2n-10)>$ 0 .
@@
Suppose there exists $(\nu,\tau)\prec(\lambda,\mu)$ such that $(\nu,\tau)$ is an arm and leg. First suppose $(\lambda,\mu)=([n-1-i,1],[1^{i}])$ for some $21\leq i\leq n-$ 2. Then
@@
where the second term on the right is zero if $i=n-2$. The dimensions of the modules are
@@
One has $([n-1-i],[1^{i}])=\beta_{n-1,i}$, and $\mathfrak{r}_{\beta_{n-1,i}}\cong\mathfrak{r}_{\beta_{n-1,1}}\cong\mathfrak{sl}(n-1)$ by the induction hypothesis, while for the other $(\nu,\tau)\prec(\lambda,\mu)$ one has $\mathfrak{r}_{(\nu,\tau)}\cong\mathfrak{sl}(S^{(\nu,\tau)})$. For $31\leq i\leq n-$ 3, this implies that

### v2

@@ 4.3.3. Arm and leg (A&L) bipartitions
This is seen via the same line of reasoning as given in the proof of [Marin:2007, Lemme 12], applied now to the natural representation $S^{\beta}$ for $\mathcal{B}_{n}$. In this case one must also consider the action of the generator $t_{1}$, for which one checks that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(t_{1})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(t_{1})+(d-1)\cdot\operatorname{Id}$. This is the same relation as for the generators $s_{1},\ldots,s_{n-1}$, so the reasoning in the proof of [Marin:2007, Lemme 12] applies. ∎
@@
Let $\rho_{\operatorname{Grp}}^{\wedge^{d}}:\mathfrak{b}_{n}\to\operatorname{End}(\Lambda^{d}(S^{\beta}))$ and $\rho_{\operatorname{Lie}}^{\wedge^{d}}:\mathfrak{b}_{n}\to\operatorname{End}(\Lambda^{d}(S^{\beta}))$ be the $\mathfrak{b}_{n}$-module structure maps for $\operatorname{Res}^{\mathbb{C}\mathcal{B}_{n}}_{\mathfrak{b}_{n}}(\Lambda^{d}(S^{\beta})\otimes\varepsilon^{\prime})$ and $\Lambda^{d}(\operatorname{Res}^{\mathbb{C}\mathcal{B}_{n}}_{\mathfrak{b}_{n}}(S^{\beta}\otimes\varepsilon^{\prime}))$, respectively. The proof again mirrors that of [Marin:2007, Lemme 12]. In this case one sees that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(t_{1})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(t_{1})-(d-1)\cdot\operatorname{Id}$, while for $s_{1},\ldots,s_{n-1}$ one again gets $\rho_{\operatorname{Lie}}^{\wedge^{d}}(s_{i})=\rho_{\operatorname{Grp}}^{\wedge^{d}}(s_{i})+(d-1)\cdot\operatorname{Id}$. Thus, for each element $s$ in a set of generators for $\mathfrak{b}_{n}$, one sees that $\rho_{\operatorname{Lie}}^{\wedge^{d}}(s)$ and $\rho_{\operatorname{Grp}}^{\wedge^{d}}(s)$ differ by a central element of $\operatorname{End}(\Lambda^{d}(S^{\beta}))$. This implies that the restrictions of $\rho_{\operatorname{Lie}}^{\wedge^{d}}$ and $\rho_{\operatorname{Grp}}^{\wedge^{d}}$ to $\mathfrak{b}_{n}^{\prime}$ are the same. ∎
@@ 4.3.4. Improper bipartitions
The structure of $\mathfrak{s}_{\lambda}$ was determined by Marin in [Marin:2007]. We summarize his results. Say that a partition $\lambda$ is a hook if $\lambda=[n-d,1^{d}]$ for some $0\leq d\leq n-1$. [^3] Marin refers to partitions of this form as équerres. Let
$E_{n}$ $=\left\{\lambda\vdash n:\text{$\lambda$ is not a hook and $\lambda\neq\lambda^{*}$}\right\},$
$F_{n}$ $=\left\{\lambda\vdash n:\text{$\lambda$ is not a hook and $\lambda=\lambda^{*}$}\right\}.$
@@
  • Given $\lambda\in F_{n}$, let $b(\lambda)$ denote the length of the diagonal in the Young diagram of $\lambda$. Then there exists a nondegenerate bilinear form [^4] Up to a nonzero scalar multiple, this is the form $(-|-)_{\left\{\lambda,\emptyset\right\}}$ on $S^{\left\{\lambda,\emptyset\right\}}=S^{\lambda}$ defined prior to Lemma 4.3.2. (|)(\;|\;) on $S^{\lambda}$, which is symmetric (resp. anti-symmetric) if $(n-b(\lambda))/2$ is even (resp. odd), and with respect to which
@@
  • For $0\leq d\leq n-1$, set $\alpha_{d}=\alpha_{n,d}=[n-d,1^{d}]$, and set $\alpha=\alpha_{n,1}$. Then $S^{\alpha_{d}}\cong\Lambda^{d}(S^{\alpha})$ as $\mathfrak{s}_{n}^{\prime}$-modules by [Marin:2007, Lemme 12]. The map $\rho_{\alpha_{d}}:\mathfrak{s}_{n}^{\prime}\to\mathfrak{sl}(\alpha_{d})$ factors through $\rho_{\alpha}:\mathfrak{s}_{n}^{\prime}\to\mathfrak{sl}(\alpha)$, and $\mathfrak{s}_{\alpha}=\mathfrak{sl}(\alpha)\cong\mathfrak{sl}(n-1)$. Then $\mathfrak{s}_{\alpha_{d}}\cong\mathfrak{s}_{\alpha}$ if $1\leq d\leq n-2$, and $\mathfrak{s}_{\alpha_{d}}=0$ otherwise.
Theorem 4.3.10 ([Marin:2007, Theorem A]).
@@ 4.3.5. Summary
$E(n)$ $=\left\{(\lambda,\mu)\in\mathcal{BP}(n)^{\circ}:\text{$(\lambda,\mu)$ is not an A\&L and $(\lambda,\mu)\neq(\mu^{*},\lambda^{*})$}\right\},$
$F(n)$ $=\left\{(\lambda,\mu)\in\mathcal{BP}(n)^{\circ}:\text{$(\lambda,\mu)$ is not an A\&L and $(\lambda,\mu)=(\mu^{*},\lambda^{*})$}\right\},$
$E\{n\}$ $=\left\{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}^{\circ}:\text{$\left\{\lambda,\mu\right\}$ is not an A\&L and $\left\{\lambda,\mu\right\}\neq\left\{\lambda^{*},\mu^{*}\right\}$}\right\},$
$F\{n\}$ $=\left\{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}^{\circ}:\text{$\left\{\lambda,\mu\right\}$ is not an A\&L and $\left\{\lambda,\mu\right\}=\left\{\lambda^{*},\mu^{*}\right\}$}\right\}.$
@@ 5.1. Statement of the main theorem
By Lemma 4.2.3 and Proposition 4.3.11, the dimensions of $\mathfrak{b}_{n}^{\prime}\subseteq\mathfrak{b}_{n}$ and $\mathfrak{d}_{n}^{\prime}\subseteq\mathfrak{d}_{n}$ for $n\in\left\{2,3,4,5\right\}$ are less than or equal to the values specified in Table 5.1. One can then verify using GAP [GAP4] that the dimensions are equal to these upper bounds, and hence the injective maps in Proposition 4.3.11 must be isomorphisms by dimension comparison. A copy of the GAP code we used to carry out these calculations is included as an ancillary file with the version of this paper posted on the arXiv (arXiv:2506.01198). ∎
$n$ $\dim(\mathfrak{b}_{n}^{\prime})\leq$ $\dim(\mathfrak{b}_{n})\leq$ $\dim(\mathfrak{d}_{n}^{\prime})\leq$ $\dim(\mathfrak{d}_{n})\leq$
2 3 5 0 1
3 19 21 11 12
4 139 141 78 79
5 1630 1632 840 841
Table 5.1. Upper bounds on dimensions of $\mathfrak{b}_{n}^{\prime}$ and $\mathfrak{b}_{n}^{\prime}$.
For the induction step we let $n\geq 6$ be fixed and we assume that Theorem 5.1.1 is true for the value n1n-1. Under this assumption we first show in Sections 5.25.9 that parts (2)–(9) of Theorem 5.1.1 are true for the value $n$; part (1) is true for all $n\geq 2$ by the identifications in (4.3.14) and by Theorem 4.3.10. Then in Section 5.10 we show that the maps (4.3.16) and (4.3.17) are surjections. Since the maps are injections by construction, this finishes the proof.
In broad strokes, the proofs of parts (2)–(9) are all the same: In addition to the given unknown Lie algebra $\mathfrak{g}$ (either $\mathfrak{d}_{\beta}$, $\mathfrak{b}_{\beta}$, $\mathfrak{b}_{\gamma}$, $\mathfrak{b}_{(\lambda,\mu)}$, etc.), we consider a semisimple subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ arising from the corresponding Weyl group of rank n1n-1. From the restriction formulas in Sections 2.5 and 3.5, we first deduce that $\mathfrak{g}$ is simple using [Marin:2007, Lemme 15]. Then in most cases, we see with the help of [Marin:2007, Lemme 13] and [Marin:2007, Lemme 14] that the rank of $\mathfrak{h}$ is sufficiently large to force $\mathfrak{g}$ to have the claimed value. There is however one case where this strategy is not entirely sufficient to determine $\mathfrak{g}$: For part (7) we must apply more detailed information about the module $S^{\left\{\lambda,\pm\right\}}$ to rule out the possibility that $\mathfrak{g}$ is either an orthogonal or symplectic Lie algebra.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (2) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The dimensions of the modules are $\dim(S^{\beta})=n$, $\dim(S^{\left\{[n-2],[1]\right\}})=n-1$, and $\dim(S^{\left\{[n-1],\emptyset\right\}})=1$. Let $\mathfrak{h}=\rho_{\beta}(\mathfrak{d}_{n-1}^{\prime})$. Then $\mathfrak{h}\subseteq\mathfrak{d}_{\beta}$ are two semisimple subalgebras of $\mathfrak{sl}(\beta)$, and from the induction hypothesis one gets $\mathfrak{h}=\mathfrak{sl}\{[n-2],[1]\}$. In particular, $\mathfrak{h}$ is semisimple of rank $\operatorname{rk}(\mathfrak{h})=n-2$, while $\operatorname{rk}(\mathfrak{d}_{\beta})\leq\operatorname{rk}(\mathfrak{sl}(\beta))=n-1$, so $\operatorname{rk}(\mathfrak{d}_{\beta})<2\cdot\operatorname{rk}(\mathfrak{h})$. Then $\mathfrak{d}_{\beta}$ is simple by [Marin:2007, Lemme 15]. Now $2\cdot\operatorname{rk}(\mathfrak{d}_{\beta})\geq 2\cdot\operatorname{rk}(\mathfrak{h})=2(n-2)>n=\dim(\beta)$, so $\mathfrak{d}_{\beta}=\mathfrak{sl}(\beta)$ by [Marin:2007, Lemme 13].
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (3) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.3.1 that the right side of (5.3.1) is a summand of the right side of (4.3.16) for the value n1n-1. This implies by induction that (5.3.1) is a surjection and $\mathfrak{r}_{(\nu,\lambda^{*})}\cong\mathfrak{sl}(\nu,\lambda^{*})$ for each $\nu\prec\lambda$. By the discussion of Section 4.3.1, $S^{(\nu,\lambda^{*})}$ and $S^{(\lambda,\nu^{*})}$ are dual as $\mathfrak{r}^{\prime}$-modules. Since $\dim(S^{(\nu,\lambda^{*})})\geq n-2\geq 4$ by Lemmas 5.3.1 and 2.2.1, and since the natural module for $\mathfrak{sl}(k)$ is self-dual only if $k=2$, it follows that $S^{(\nu,\lambda^{*})}\not\cong S^{(\lambda,\nu^{*})}$ as $\mathfrak{r}^{\prime}$-modules.
@@
Then $\mathfrak{b}_{(\lambda,\lambda^{*})}$ is a simple Lie algebra by [Marin:2007, Lemme 15]. Since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})$, the preceding inequalities also imply that $2\cdot\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})\leq\dim(S^{(\lambda,\lambda^{*})})<4\cdot\operatorname{rk}(\mathfrak{b}_{(\lambda,\lambda^{*})})$. Now since $\dim(S^{(\lambda,\lambda^{*})})\geq\binom{n}{n/2}$, it follows from [Marin:2007, Lemme 14] that $\mathfrak{b}_{(\lambda,\lambda^{*})}=\mathfrak{osp}(\lambda,\lambda^{*})$. ∎
@@
    for some $1\leq i\leq n-2$, and $(\nu,\tau)$ is the unique arm and leg such that $(\nu,\tau)\prec(\lambda,\mu)$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (4) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.4.1 that the right side of (5.4.2) is a summand of the right side of (4.3.16) for the value n1n-1, except perhaps that a summand of the form $\mathfrak{r}_{\beta_{n-1,1}}$ or $\mathfrak{r}_{\gamma_{n-1,1}}$ may be replaced by an isomorphic summand of the form $\mathfrak{r}_{\beta_{n-1,d}}$ or $\mathfrak{r}_{\gamma_{n-1,d}}$ for some $1\leq d\leq n-2$, and a summand of the form $\mathfrak{s}_{\alpha_{n-1,1}}$ may be replaced by an isomorphic summand $\mathfrak{s}_{\alpha_{n-1,d}}$ for some $1\leq d\leq n-3$. This implies by the induction hypothesis that (5.4.2) is a surjection, and implies for each $(\nu,\tau)\prec(\lambda,\mu)$ that $\mathfrak{r}_{(\nu,\tau)}$ is a simple Lie algebra acting irreducibly on $S^{(\nu,\tau)}$.
Let $\mathfrak{h}=\rho_{(\lambda,\mu)}(\mathfrak{r}^{\prime})$. Then $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{b}_{(\lambda,\mu)}$ such that the restriction of $S^{(\lambda,\mu)}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. Since $\mathfrak{h}\subseteq\mathfrak{b}_{(\lambda,\mu)}\subseteq\mathfrak{sl}(\lambda,\mu)$, then $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{b}_{(\lambda,\mu)})\leq\dim(S^{(\lambda,\mu)})$. Our goal now, as in the proof of Lemma 5.2.1, is to show that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{(\lambda,\mu)})$. This will imply first by [Marin:2007, Lemme 15] that $\mathfrak{b}_{(\lambda,\mu)}$ is a simple Lie algebra, and then by [Marin:2007, Lemme 13] that $\mathfrak{b}_{(\lambda,\mu)}=\mathfrak{sl}(\lambda,\mu)$.
@@
Suppose $\mu=[1]$ and $\lambda$ is a hook partition. Then $\lambda=\alpha_{n-1,d}=[n-1-d,1^{d}]$ for some $1\leq d\leq n-3$; the values $d=0$ and d=n2d=n-2 are excluded by the assumption that $(\lambda,\mu)$ is not an arm and leg. In this case, (5.4.1) becomes
@@
From Section 4.3.4, $\mathfrak{r}_{(\alpha_{n-1,d},\emptyset)}\cong\mathfrak{sl}(\alpha_{n-1,1})\cong\mathfrak{sl}(n-2)$. For $d\neq n-3$, the bipartition $(\alpha_{n-2,d},[1])$ is not an arm and leg, and hence $\mathfrak{r}_{(\alpha_{n-2,d},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d},[1])})$, while for d=n3d=n-3, one has $(\alpha_{n-2,d},[1])=([1^{n-2}],[1])=\gamma_{n-1,n-2}$. By induction, $\mathfrak{r}_{\gamma_{n-1,n-2}}\cong\mathfrak{sl}(\gamma_{n-1,1})\cong\mathfrak{sl}(n-1)$, so by dimension comparison the isomorphism $\mathfrak{r}_{(\alpha_{n-2,d},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d},[1])})$ is true for d=n3d=n-3. Similarly, one sees for all $1\leq d\leq n-3$ that $\mathfrak{r}_{(\alpha_{n-2,d-1},[1])}\cong\mathfrak{sl}(S^{(\alpha_{n-2,d-1},[1])})$. Then
@@
$2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=(n-2)\cdot\binom{n-2}{d}+(2n-10)>0.$
@@
Suppose there exists $(\nu,\tau)\prec(\lambda,\mu)$ such that $(\nu,\tau)$ is an arm and leg. First suppose $(\lambda,\mu)=([n-1-i,1],[1^{i}])$ for some $1\leq i\leq n-2$. Then
@@
where the second term on the right is zero if i=n2i=n-2. The dimensions of the modules are
@@
One has $([n-1-i],[1^{i}])=\beta_{n-1,i}$, and $\mathfrak{r}_{\beta_{n-1,i}}\cong\mathfrak{r}_{\beta_{n-1,1}}\cong\mathfrak{sl}(n-1)$ by the induction hypothesis, while for the other $(\nu,\tau)\prec(\lambda,\mu)$ one has $\mathfrak{r}_{(\nu,\tau)}\cong\mathfrak{sl}(S^{(\nu,\tau)})$. For $1\leq i\leq n-3$, this implies that